OpenStudy (anonymous):

calculate the limit as t-->0 for (sqr(1+t)-1)/t

OpenStudy (turingtest):

multiply by \[\frac{\sqrt{1+t}+1}{\sqrt{1+t}+1}\]

OpenStudy (anonymous):

yeah just realized that

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you da bomb

OpenStudy (anonymous):

Woops, lol, I was about to suggest l'Hopital's rule. That's a better answer.

OpenStudy (anonymous):

hah

OpenStudy (anonymous):

no worries